OpenStudy (anonymous):

help ...if (4k-1)/2=k,then k=.....

OpenStudy (bahrom7893):

multiply both sides by 2: 4k-1 = 2k add 1 to both sides: 4k=2k+1 subtract 2k from both sides: 2k = 1 divide both sides by 2: k = 1/2

sam (.sam.):

k=(1)/(2)

OpenStudy (anonymous):

k= 1/2

sam (.sam.):

(4k-1)/(2)=k 4k-1=2k 2k-1=0 2k=1 k=(1)/(2)